OpenStudy (anonymous):

A can of soup is in the shape of a cylinder. The diameter of the can is 8 centimeters, and the height is 10 centimeters. What is the surface area of the can, rounded to the nearest hundredth square centimeter? (Use 3.14 for π.) 351.68 cm 2 251.2 cm 2 502.4 cm 2 100.48 cm 2

OpenStudy (anonymous):

what have you tried

OpenStudy (anonymous):

I know that B is wrong

OpenStudy (anonymous):

I'm using meter units (habit). Area of top is\[\pi r ^{2} = \pi (0.04)^2 = 0.00502654824574\]Area of bottom is the same. \[0.00502654824574\]Area of the vertical component is\[height \times length = height \times 2 \pi r = 0.1 \times 2 \pi (0.04)=0.02513274122872\]Add them up gives \[0.0351858377202m^{2}\] or in units of CM squared (100x100) is \[351.858377202cm^{2}\]